# MOOSE-Chem: Large Language Models for Rediscovering Unseen Chemistry Scientific Hypotheses

### v3

@@ 5.3 How Well can LLMs perform $R(h)$?
Table 8 shows the relation between the #Matched $i$ and average ranking ratio (the lower, the better). It shows a clear trend that the more ground truth $i$ is leveraged, the better ranking score $h$ can have. It indicates that that $h$ with a higher ranking ratio are more likely to be matched with better $i$.
Table 9 shows the relation between the GPT-4o evaluated MS and the average ranking ratio. For $h$ with a positive MS, there is a trend that the higher the MS, the better the average rank ratio (if MS $\in$ [2,4]). However, the disadvantage of those $h$ without a positive MS is not very significant. It seems that LLMs have a certain ability to rank good $h$ higher. But it is not sure how significant it is, because a part of the reason of this results is that those $h$ generated without groundtruth $i$ could be also in high quality.
@@ 6.2 Results
   w/o EU 3.765 2.730
   w/o multi-step 3.588 2.452
Table 10: Experiments and ablation study. The Matched Score is evaluated by GPT-4o.
@@
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ the top 4 generated $h$ with the highest MS by GPT-4o are evaluated again by two experts in terms of MS. Table 11 shows the expert evaluation results. Here the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the chemistry hypothesis, which is published in Nature, Science or a similar level.
@@
7 Case Study
Here, we show a case study of a generated $h$ (from MOOSE-Chem with $|I|$=300), corresponding ground truth $h$, and an expert’s analysis (gives a MS as 3). More case studies can be seen in § A.8.
Ground truth $h$: The main hypothesis is that a nitrogen-doped ruthenium (Ru) electrode can effectively catalyze the reductive deuteration of (hetero)arenes in the presence of $\mathbf{D_{2}O}$, leading to high deuterium incorporation into the resulting saturated cyclic compounds. The findings validate this hypothesis by demonstrating that this electrocatalytic method is highly efficient, scalable, and versatile, suitable for a wide range of substrates.
@@ 8 Conclusion
We investigate this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We propose a fundamental assumption to break up this seemingly-impossible-to-solve central question into three smaller, more practical and executable fundamental questions. Then, we investigate LLM’s ability on each of them. To do it, we construct a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also develop an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments show that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@ Acknowledgments
This research/project is supported by the Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005).
@@ A.1 Full Proof / Derivation for the Fundamental Assumption
We denote the full (chemistry) literature as $I$, such that $P(I)=1$. Then we can have:
@@
Here, $P(h|b,i_{1},\ldots,i_{k})$ is used to obtain $h$, and $P(i_{j+1}|b,i_{1},\ldots,i_{j},I)$ is used to obtain each $i$. However, they do not precisely reflect how chemistry researchers do, and might not be practicable. One of the main reasons is that researchers tend to think small step by small step. It would be very challenging to think in terms of a big step without breaking it into several small steps.
@@
As a result, to achieve the final $h_{k}$, we need to obtain $h_{1},\ldots,h_{k-1}$ first (if $k>1$). In addition, seeing $h$ as a “state”, and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}|b,h_{k-1},i_{k})$ and $P(i_{k+1}|b,h_{k},I)$ correspondingly indicates a property very similar to the Markov property: (1) a new $h$ only depends on $b$, its previous $h$, and the current $i$; and (2) an $i$ only depends on $b$, $I$, and the current $h$.
@@
Although starting from $k>1$, Derivation 18 covers the situation when $k=1$ in Equation 8.
@@
A.2 The Full Instruction for Benchmark Checking
@@ A.3 Prompt to obtain $R(h)$
A.4 Related Works on Reasoning
@@
Another relevant reasoning type is commonsense reasoning (Yang et al., 2020; 2023a). Scientific discovery can be seen as an opposite task, which is to reason far outside of commonsense, even to discover the unknown knowledge.
A.5 Prompt to GPT-4o for Matched Score
@@
A.6 Generated Hypotheses With Low Matched Score Are Not Necessarily Bad
MS only measures the similarity between the generated $h$ and the ground truth $h$. Receiving a MS as 0 or 1 does not mean the generated $h$ is in bad. Only real lab experiments can check each $h$.
A.7 Agreement Between Expert Evaluation and GPT-4o Evaluation
#Comparison Examples Hard Consistency Score Soft Consistency Score
@@ A.7 Agreement Between Expert Evaluation and GPT-4o Evaluation
Table 12 shows the agreement between expert evaluation and automatic evaluation (by GPT-4o) on MS. Hard consistency is assigned to 1 only if the two scores are exact the same, else is assigned to 0. Soft consistency is assigned to 1 only if the absolute difference between the two scores is less than 2, else is assigned to 0.
@@
A.8 More Case Study
A.8.1 Case 1 (MS by Expert: 4)

### v4

@@ 5.3 How Well can LLMs perform $R(h)$?
Table 8 shows the relation between the #Matched $i$ and average ranking ratio (the lower, the better). It shows a clear trend that the more ground truth $i$ is leveraged, the better ranking score $h$ can have. It indicates that $h$ with a higher ranking ratio is more likely to be matched with better $i$.
Table 9 shows the relation between the GPT-4o evaluated MS and the average ranking ratio. There is a trend that the higher the MS, the better the average rank ratio (when MS $\in$ [2,4]). However, the disadvantage of those $h$ without a positive MS is not very significant. It seems that LLMs have a certain ability to rank good $h$ higher. But it is not sure how significant it is, because a part of the reason for these results is that those $h$ generated without groundtruth $i$ could be also in high quality.
@@ 6.2 Results
   w/o multi-step 3.765 2.730
   w/o multi-step & EU 2.863 2.578
Table 10: Experiments and ablation study. The Matched Score (MS) is evaluated by GPT-4o (this table), Claude-3.5-Sonnet (Table 13), and Gemini-1.5-Pro (Table 14).
@@
To better understand the performance of MOOSE-Chem in this real copilot setting, for each $b$ the top 4 generated $h$ with the highest MS by GPT-4o are evaluated again by two experts in terms of MS. Table 11 shows the expert evaluation results. Here the top MS is the highest MS for each $b$, out of the 4 expert evaluated $h$ for this $b$. Note that MS rated as three is already very high. Illustrated in Table 6, it means the generated $h$ by MOOSE-Chem (that has not seen $h$) in the real copilot setting covers two main innovations of the chemistry hypothesis, which is published in Nature, Science or a similar level. Some case studies can be seen in § A.16.
@@ 7 Conclusion
We investigate this central question: “Can LLMs automatically discover novel and valid chemistry (including material science) research hypotheses (even those which deserve a publication in Nature, Science, or a similar level) given only a chemistry research background (consisting of a research question and/or a background survey), without limitation on the domain of the research question?”. We propose a fundamental assumption to break up this seemingly impossible-to-solve central question into three smaller, more practical, and executable fundamental questions. Then, we investigate LLM’s ability on each of them. To do it, we construct a benchmark consisting of chemistry and material science papers published and only be public in 2024. We also develop an LLM-based multi-agent framework consisting of three stages reflecting the three smaller fundamental questions. Experiments show that the framework (runs in a copilot in-the-wild setting, with LLMs with training data up to October 2023) can rediscover many hypotheses with very high similarity with the ground truth ones, covering the main innovations.
@@ Acknowledgments
This work is supported by the Shanghai Municipal Science and Technology Major Project. This work is supported by Shanghai Artificial Intelligence Laboratory. This research/project is supported by the Ministry of Education, Singapore under its MOE Academic Research Fund Tier 2 (STEM RIE2025 Award MOE-T2EP20123-0005).
@@ A.1 Full Proof / Derivation for the Fundamental Assumption
Our goal is to transform the seemingly impossible-to-solve $P(h|b)$ into an equivalent form, where each step in the equivalent form is practical and executable. Denoting the full (chemistry) literature as $I$, such that $P(I)=1$. Then a straightforward way of decomposing $P(h|b)$ is by the chain rule based on Equation 4:
@@
Equation 7 describes the process of $P(h|b)$ in terms of the knowledge-searching perspective. However, $P(h|b,i_{1},\ldots,i_{k})$ and $P(i_{j}|b,i_{1},\ldots,i_{j-1},I)$ might not be enough practicable, and do not precisely reflect how chemistry researchers find a new $i$. One of the main reasons is that researchers tend to think small step by small step. It would be very challenging to think in terms of a big step without breaking it into several small steps.
@@
As a result, to achieve the final $h_{k}$, we need to obtain $h_{1},\ldots,h_{k-1}$ first (if $k>1$). In addition, seeing $h$ as a “state”, and $i$ as an “action”, obtaining $h$ and $i$ through $P(h_{k}|b,h_{k-1},i_{k})$ and $P(i_{k+1}|b,h_{k},I)$ correspondingly indicates a Markov property: (1) a new $h$ only depends on $b$, its previous $h$, and the current $i$; and (2) an $i$ only depends on $b$, $I$, and the previous $h$.
@@
Although starting from $k>1$, Derivation 17 covers the situation when $k=1$ in Equation 7.
@@
A.2 MOOSE-Chem Overview I/O Figure
Figure 2 shows the input and output overview of the MOOSE-Chem framework.
A.3 The Full Instruction for Benchmark Checking
@@ A.4 Prompt to obtain $R(h)$
A.5 Automatic Evaluation by Claude and Gemini
5 4 3 2 1 0 Total
w/ background survey
Average MS (GPT-4o) 2 9 18 17 5 0 51
Average MS (Claude-3.5-Sonnet) 4 19 15 10 3 0 51
Average MS (Gemini-1.5-Pro) 2 13 17 8 11 0 51
Top MS (GPT-4o) 28 1 19 3 0 0 51
Top MS (Claude-3.5-Sonnet) 33 7 10 1 0 0 51
Top MS (Gemini-1.5-Pro) 20 18 0 12 1 0 51
Top MS (Experts) 9 12 22 6 2 0 51
w/o background survey
Average MS (GPT-4o) 1 7 17 19 7 0 51
Average MS (Claude-3.5-Sonnet) 7 24 18 2 0 0 51
Average MS (Gemini-1.5-Pro) 4 9 14 15 5 4 51
Top MS (GPT-4o) 25 2 19 5 0 0 51
Top MS (Claude-3.5-Sonnet) 31 19 1 0 0 0 51
Top MS (Gemini-1.5-Pro) 19 19 1 11 0 1 51
To investigate whether the results and corresponding conclusions in the main text are caused by the usage of GPT-4o for automatic evaluation, here we use Claude-3.5-Sonnet and Gemini-1.5-Pro to evaluate all of the results that have been evaluated by GPT-4o.
Table 12 covers the contents in Table 7, but with more results on using Claude-3.5-Sonnet and Gemini-1.5-Pro for automatic evaluation. When using different LLMs for automatic evaluation, the instruction is the same (can be found in § A.12). The robust results indicate again that LLMs are quite capable of associating known knowledge into unknown knowledge that has a high probability to be valid (very close to oh).
Method Top MS Average MS
SciMON (Wang et al., 2024b) 3.824 3.529
MOOSE (Yang et al., 2024a) 3.902 3.559
Qi et al. (2024) 3.431 3.092
MOOSE-Chem 4.471 3.697
   w/o multi-step 4.216 3.592
   w/o multi-step & EU 3.941 3.614
Table 13: Experiments and ablation study. The Matched Score is evaluated by Claude-3.5-Sonnet.
Method Top MS Average MS
SciMON (Wang et al., 2024b) 2.980 2.618
MOOSE (Yang et al., 2024a) 3.039 2.690
Qi et al. (2024) 2.216 1.846
MOOSE-Chem 3.686 2.443
   w/o multi-step 3.588 2.529
   w/o multi-step & EU 2.902 2.631
Table 14: Experiments and ablation study. The Matched Score is evaluated by Gemini-1.5-Pro.
Table 13 and Table 14 evaluate the same hypotheses with Table 10, but using Claude-3.5-Sonnet and Gemini-1.5-Pro for automatic evaluation correspondingly (instead of GPT-4o). The results indicate the robustness of MOOSE-Chem and its components.
A.6 More Analysis on EU
MS threshold only non-EU branch only EU branches only EU-recombination branch
5 16 46 20
4 19 54 24
Table 15: Number of hypotheses receiving high Matched Score (MS) from only non-EU branch, only EU branches, and only EU-recombination branch. Only the hypotheses with a MS that is higher than the MS threshold are counted.
Table 15 shows the number of hypotheses receiving high Matched Score from only non-EU branch, only EU branches, and only EU-recombination branch. Here only non-EU branch can be seen as the hypotheses obtained directly without mutations. The hypotheses are from the same experiment in Table 10.
The result indicates that about one-third of high-quality hypotheses can be obtained directly without mutations. In addition, the recombination branch contains more high-quality hypotheses than the only non-EU branch.
A.7 Effect of Significance Feedback
5 4 3 2 1 0
w/ significance feedback
Average MS 4 19 15 10 3 0
Top MS 33 7 10 1 0 0
w/o significance feedback
Average MS 8 28 11 3 1 0
Top MS 34 13 4 0 0 0
Table 16: Effect of significance feedback (evaluated by Claude-3.5-Sonnet).
Table 16 presents an ablation study on the significance feedback. The results with significance feedback are from Table 12.
The results indicate that not using significance feedback can even lead to a better performance in terms of the Matched Score metric. We attribute this phenomenon to LLM’s ability on creativity: when asked to generate significant hypotheses, LLMs tend to be more deviate from the existing information for more possible significance, resulting in a lower matched score. However, we should note that the matched score only measures the matching degree of one given groundtruth hypothesis, and it is possible that the more deviated one is more significant.
A.8 Ranking of Groundtruth Hypotheses
Overall Validness Novelty Significance Potential
Average Rank Ratio 0.65 0.75 0.76 0.73 0.70
Table 17: Average rank ratio ($\downarrow$) of the ground truth hypotheses (mixed with generated hypotheses)
Intuitively if we rank the original hypothesis with the generated hypothesis, the original hypothesis may be ranked at the top for most of the time. But is it?
Table 17 shows the result, where we assign each ground truth hypothesis with a reward value $R(h)$ (in terms of validness, novelty, significance, and potential), and calculate its average rank ratio regarding the framework-generated hypotheses.
Surprisingly, the ground truth hypotheses are not ranked to the top. There are three possible reasons:
  1. 1.
    LLM does poorly on ranking hypotheses;
  2. 2.
    The generated hypotheses tend to describe their novelty and significance (although they are prompted to not to), which might influence the judgment;
  3. 3.
    The generated hypotheses may surpass the original in quality.
  4. 4.
    The generated hypotheses may sometimes have more details than the groundtruth one (since the iterative usage of clarity feedback and refinement).
A.9 Influence of Research Background Options to Inspiration Retrieval
Strict Background Background Survey Hit Ratio (top 20%) Hit Ratio (top 4%) Hit Ratio (top 0.8%)
96.7% 83.7% 60.8%
95.1% 77.8% 54.2%
96.7% 80.1% 57.8%
A.10 Discussion on Hallucination and Scientific Discovery
In contrast to the traditional understanding that hallucination is purely a bad thing, LLM’s scientific discovery ability in fact counts on its hallucination ability to find novel hypotheses: a novel hypothesis would not have been observed by itself, therefore all novel hypotheses come from the class of hallucination.
In essence, the research development of LLMs for automated scientific hypothesis discovery is to develop how to better leverage LLMs to hallucinate an unseen hypothesis that has more possibility to be valid.
A.11 Other Related Works
A.11.1 Reasoning
@@
Another relevant reasoning type is commonsense reasoning (Yang et al., 2020; 2023a). Scientific discovery can be seen as an opposite task, which is to reason far outside of commonsense, even to discover unknown knowledge.
A.11.2 Retrieval
The retrieval of inspiration is a retrieval task, and RAG (Lewis et al., 2020) also works on retrieval. The main difference is that the current RAG method would most likely retrieve the information that is semantically the most similar to the input information (research background), while here our goal is to retrieve those information that was not known to be related to the input information before, but in fact is related. We assume that LLMs might have the ability to do it.
A.11.3 Self Consistency
Self-consistency (Wang et al., 2023; Chen et al., 2023) might have a similar looking to the “evolutionary unit” (EU), as they all have expansion to several branches, and finally collect these branches into one.
A key difference is that EU is to explore more diverse options to choose the optimal one, while self-consistency is to find consistent voting between options.
A.12 Prompt to GPT-4o for Matched Score
@@
A.13 Generated Hypotheses With Low Matched Score Are Not Necessarily Bad
MS only measures the similarity between the generated $h$ and the ground truth $h$. Receiving an MS as 0 or 1 does not mean the generated $h$ is bad. Only real lab experiments can check each $h$.
A.14 Evaluation Agreement Between Expert Evaluation and GPT-4o Evaluation
#Comparison Pairs Hard Consistency Score Soft Consistency Score
@@ A.14 Evaluation Agreement Between Expert Evaluation and GPT-4o Evaluation
#Comparison Pairs Hard Consistency Score Soft Consistency Score
48 0.438 0.854
Table 20: Consistency score between experts in expert evaluation.
Table 19 shows the agreement between expert evaluation and automatic evaluation (by GPT-4o) on MS. Hard consistency is assigned to 1 only if the two scores are exactly the same, else is assigned to 0. Soft consistency is assigned to 1 only if the absolute difference between the two scores is less than 2, else is assigned to 0.
@@
A.15 Agreement Between Expert Evaluation
Table 20 shows the consistency score between experts. It is conducted by having the third chemistry expert check the randomly sampled expert evaluation from the previous two chemistry experts. Hard consistency is assigned to 1 only if the two scores are exactly the same, else is assigned to 0. Soft consistency is assigned to 1 only if the absolute difference between the two scores is less than 2, else is assigned to 0.
The results show a high consistency between experts, indicating a high reliability of reference-based expert evaluation.
A.16 Case Studies
A.16.1 Case 1 (MS by Expert: 3)
Ground truth $h$ (Bu et al., 2024): The main hypothesis is that a nitrogen-doped ruthenium (Ru) electrode can effectively catalyze the reductive deuteration of (hetero)arenes in the presence of $\mathbf{D_{2}O}$, leading to high deuterium incorporation into the resulting saturated cyclic compounds. The findings validate this hypothesis by demonstrating that this electrocatalytic method is highly efficient, scalable, and versatile, suitable for a wide range of substrates.
A.16.2 Case 2 (MS by Expert: 4)